# T-CBF: Traversability-based Control Barrier Function to Navigate Vertically Challenging Terrain

### v1

@@ Abstract
Safety has been of paramount importance in motion planning and control techniques and is an active area of research in the past few years. Most safety research for mobile robots target at maintaining safety with the notion of collision avoidance. However, safety goes beyond just avoiding collisions, especially when robots have to navigate unstructured, vertically challenging, off-road terrain, where vehicle rollover and immobilization is as critical as collisions. In this work, we introduce a novel Traversability-based Control Barrier Function (T-CBF), in which we use neural Control Barrier Functions (CBFs) to achieve safety beyond collision avoidance on unstructured vertically challenging terrain by reasoning about new safety aspects in terms of traversability. The neural T-CBF trained on safe and unsafe observations specific to traversability safety is then used to generate safe trajectories. Furthermore, we present experimental results in simulation and on a physical Verti-4 Wheeler (V4W) platform, demonstrating that T-CBF can provide traversability safety while reaching the goal position. T-CBF planner outperforms previously developed planners by 30% in terms of keeping the robot safe and mobile when navigating on real world vertically challenging terrain.
@@ I Introduction
Field robots deployed in real-world applications often need to traverse unstructured, complex, and unpredictable off-road terrain, where safety becomes a top priority. The robotics community has developed numerous techniques to address motion planning safety: Traditional model-based control methods, including Hamilton-Jacobi Reachability Analysis [1], Control Barrier Functions [2], Control Lyapunov Functions [3], and Model Predictive Control [4], have proven effective but struggles with scalability and generalizability across diverse environments. To overcome these limitations, data-driven approaches [5, 6, 7, 8, 9] are emerging, offering improved adaptability to new environments but at a price of lacking formal safety guarantees.
@@
To address this challenge, this work proposes a Traversability-based Control Barrier Function (T-CBF) to enable traversability safety for navigation of wheeled robots on unstructured and vertically challenging terrain using onboard perception inputs. The main contributions are threefolds: (1) The introduction of a novel approach to construct a CBF that extends the concept of safety to traversability by incorporating perception-based information. (2) Demonstrating the generalizability of T-CBF to previously unseen unstructured environments. (3) A statistical evaluation and experimental verification of T-CBF on an autonomous Verti-4 Wheeler (V4W) [11] platform in diverse environments. The results demonstrate that T-CBF provides traversability safety beyond collision avoidance while navigating complex, vertically challenging, unstructured environments.
@@ II-C Learned Control Barrier Functions
With the growing adoption of data-driven methods, CBFs are becoming increasingly significant in robotics applications. In general, synthesizing a CBF requires a mathematical formulation that can either be learned by a neural network or directly incorporated as a constraint in an optimization framework. Xiao et al. [31], designed a neural network layer that is compatible with any neural CBF, enabling training based on a mathematical formulation specific to a given scenario, thereby ensuring safety guarantees. For approaches utilizing perception data, Long et al. [32] developed a CBF that is created online using range sensing for obstacle avoidance, and Dawson et al. [33] designed a hybrid controller incorporating a learned CBF to navigate using LiDAR point cloud data. Similarly, Abidi et al.  [34] leveraged camera sensor images to construct a vision-based learned CBF for obstacle avoidance while maintaining a safe following distance. Additionally, Harms et al. [6] introduced a neural CBF-based safety filter for quadrotor navigation in unknown environments. These approaches demonstrate the effectiveness of machine learning in designing scalable CBFs that generalize to new environments. Reinforcement learning has also been used in conjunction with CBFs to develop safe navigation policies for obstacle avoidance [35, 36, 37].
@@ IV-A Traversability-based Control Barrier Function
Formulating safety constraints for high dimensional systems on unstructured terrain is complicated due factors such as suspension and tire deformation, varying tire-terrain friction, vehicle weight distribution and momentum, etc. To address this, we utilize manually driven runs to collect data and identify unsafe states (Fig. 1) of the robot based on the following criteria: Denote the robot state as $\textbf{x}_{t}=(x_{t},y_{t},z_{t},r_{t},p_{t},\phi_{t})$, where the first three are translational $(x,y,z)$ and last three are rotational (roll, pitch, yaw) components respectively along x, y, and z axis, and control input as $\textbf{u}_{t}=(v_{t},\omega_{t})$, where $v$ is translational velocity and $\omega$ is angular velocity which also correlates to steering curvature.
@@
Figure 2: T-CBF Neural Network Architecture: The observations (elevation maps) are processed through a neural network as CBF and control actions are processed through an encoder. The loss function helps learn a CBF enforcing satisfaction of Eqns. (4a), (4b), and (4c). The output of the T-CBF is a safe set region (dotted red boundary) on elevation map.
@@ IV-B Learning T-CBF from Observations
$h[o(f(x_{t},u))]\geq-\alpha_{h}(h(o,x_{t})),\>\>\>\>\>\forall u\in\mathcal{U},\forall x\in\mathcal{X}$ (4c)
where $10\leq\alpha_{h}<$ 1 is an extended class $\mathcal{K}_{\infty}$ function. We formulate our CBF similar to Harms et al. [6] and use the switching set approach to avoid the observation dynamics approximation that comes from differentiating $h(o,x)$ and results in $\frac{d}{dt}h(o,x)=h[o(f(x_{t},u))]$ which represents the state dynamics of the system.
@@ V-A Robot and Testbed
Figure 4: Experimental evaluation of T-CBF. A & B: Simulation run of robot avoiding unsafe regions while navigating challengin terrain. 1: Robot initialization with start and goal pose in safe region. 2-4: Robot navigating unstructured vertically challenging terrain and avoiding nontraversable unsafe regions to reach the goal position. The bottom images show traversability-based safe regions (dotted red boundary) on elevation map computed by T-CBF and resulted trajectories in green.
@@ V-B T-CBF Validation
To evaluate the safety performance of T-CBF, we configure the testbed to simulate unsafe traversable scenarios. We conduct experiments using the WM-VCT and TAL algorithms, which lack explicit safety constraints, and compare their results with T-CBF. Throughout our experiments, we assume that the robot’s start and goal positions are located in a safe region. Unlike the other planners, T-CBF successfully prevents the robot from entering in unsafe states and getting immobilized. The robot remains mobile and capable of searching alternate paths towards goal position. In contrast, WM-VCT and TAL become immobilized while attempting to navigate the unsafe terrain. The results are presented in Fig. 3. This experiment validates that T-CBF is able to provide traversability safety when navigating unstructured, vertically challenging terrain.
@@ V-C Performance Comparison
After validating T-CBF’s traversability safety beyond collision avoidance, we compare its performance with WM-VCT and TAL in both simulation and real-world experiments.

### v2

@@ Abstract
Control Barrier Function, Unstructured Environment, Safe Navigation, Field Robotics, Traversability
Index Terms:
Control Barrier Function, Unstructured Environment, Safe Navigation, Field Robotics, Traversability
@@ I Introduction
Field robots deployed in real-world applications often need to traverse unstructured, complex, and unpredictable off-road terrain, where safety becomes a top priority. The robotics community has developed numerous techniques to address motion planning safety: Traditional model-based control methods, including Hamilton-Jacobi Reachability Analysis [1], Control Barrier Functions [2], Control Lyapunov Functions [3], and Model Predictive Control [4], have proven effective but struggles with scalability and generalizability across diverse environments. To overcome these limitations, data-driven approaches [5, 6] are emerging, offering improved adaptability to new environments but at a price of lacking formal safety guarantees.
@@
To address this challenge, this work proposes a Traversability-based Control Barrier Function (T-CBF) to enable traversability safety for navigation of wheeled robots on unstructured and vertically challenging terrain using onboard perception inputs. The main contributions are threefolds: (1) The introduction of a novel approach to construct a CBF that extends the concept of safety to traversability by incorporating perception-based information. (2) Demonstrating the generalizability of T-CBF to previously unseen unstructured environments. (3) A statistical evaluation and experimental verification of T-CBF on an autonomous Verti-4 Wheeler (V4W) [11] platform in diverse environments. The results demonstrate that T-CBF providesmay prevent the theoretical guarantees outlined in Section V from fully applying in real-world scenarios. To support the practical effectiveness of our learned certificates, the following section presents experimental results demonstrating that the controller consistently operates safely and achieves its objectives across diverse simulated and physical environments. traversability safety beyond collision avoidance while navigating complex, vertically challenging, unstructured environments.
@@ II-C Learned Control Barrier Functions
With the growing adoption of data-driven methods, CBFs are becoming increasingly significant in robotics applications. In general, synthesizing a CBF requires a mathematical formulation that can either be learned by a neural network or directly incorporated as a constraint in an optimization framework. For approaches utilizing perception data, Long et al. [32] developed a CBF that is created online using range sensing for obstacle avoidance, and Dawson et al. [33] designed a hybrid controller incorporating a learned CBF to navigate using LiDAR point cloud data. Similarly, Abidi et al.  [34] leveraged camera sensor images to construct a vision-based learned CBF for obstacle avoidance while maintaining a safe following distance. Additionally, Harms et al. [6] introduced a neural CBF-based safety filter for quadrotor navigation in unknown environments. These approaches demonstrate the effectiveness of machine learning in designing scalable CBFs that generalize to new environments. Reinforcement learning has also been used in conjunction with CBFs to develop safe navigation policies for obstacle avoidance [35, 36, 37].
@@ IV-A Traversability-based Control Barrier Function
[…]
@@
Figure 2: T-CBF Neural Network Architecture: The observations (elevation maps) are processed through a neural network as CBF and control actions are processed through an encoder. The loss function helps learn a CBF enforcing satisfaction of Eqns. (4a), (4b), and (4c). The output, safe region (dotted red boundary), is obtained by sampling observations on elevation map from a fixed position with different control commands.
@@ IV-B Learning T-CBF from Observations
$h[o(f(x_{t},u))]\geq-\alpha_{h}(h(o,x_{t})),\>\>\forall u\in\mathcal{U}_{\text{safe}},\forall x\in\mathcal{X}_{\text{safe}}$ (4c)
where $0\leq\alpha_{h}<1$ is an extended class $\mathcal{K}_{\infty}$ function. We formulate our CBF similar to Harms et al. [6] and use the switching set approach to avoid the observation dynamics approximation that comes from differentiating $h(o,x)$ and results in $\frac{d}{dt}h(o,x)=h[o(f(x_{t},u))]$ which represents the state dynamics of the system.
@@ V-A Robot and Testbed
Figure 4: Experimental evaluation of T-CBF. A & B: Simulation run of robot avoiding unsafe regions while navigating challengin terrain. 1: Robot initialization with start and goal pose in safe region. 2-4: Robot navigating unstructured vertically challenging terrain and avoiding nontraversable unsafe regions to reach the goal position with corresponding bottom images showing traversability-based safe regions (dotted red boundary) on elevation map computed by T-CBF and resulted trajectories in green.
@@ V-B T-CBF Validation
To evaluate the safety performance of T-CBF, we configure the testbed to simulate unsafe traversable scenarios. We conduct experiments using the WM-VCT and TAL algorithms, which lack explicit safety constraints, and compare their results with T-CBF. Throughout our experiments, we assume that the robot’s start and goal positions are located in a safe region. Unlike the other planners, T-CBF successfully prevents the robot from entering in unsafe states and getting immobilized. The robot remains mobile and capable of searching alternate paths towards goal position. In contrast, WM-VCT and TAL become immobilized while attempting to navigate the unsafe terrain. The results are presented in Fig. 3. This experiment validates that T-CBF is able to provide traversability safety when navigating unstructured, vertically challenging terrain. While this validation analysis provides useful evidence, it does not constitutes a formal proof. Practical limitations like perception and localization noise and learned neural network architecture may prevent the theoretical grantees in Section IV from fully applying in real world scenarios. To support the practical effectiveness of our learned certificates, the following section presents experimental results demonstrating that the controller consistently operates safely and achieves its objectives across diverse simulated and physical environments.
@@ V-C Performance Comparison
After validating T-CBF’s traversability safety beyond collision avoidance, we compare its performance with WM-VCT and TAL in both simulation and real-world experiments. Differential locks of V4W are disabled for the real-world experiments.